OpenStudy (anonymous):

List three different ways to write 5^11 as the product of two powers. Explain why all three of your expressions are equal to 5^11

OpenStudy (e.mccormick):

What problem are you having with this one?

OpenStudy (anonymous):

idk how to wright this and i don't know how to list two different products and get the same for 5^11

OpenStudy (e.mccormick):

Rules of exponents: \[x^{m}\cdot x^{n}=x^{n+n}\\ (x^{m})^n=x^{m\cdot n}\]By using these, you can change things around.

OpenStudy (anonymous):

idk how to do that i am just getting into this

OpenStudy (e.mccormick):

OK. You know what \(5^{11}\) means? What would be done to 5, right?

OpenStudy (anonymous):

yes

OpenStudy (e.mccormick):

Another way of saying 5 is \(5^1\). That works for anything. \(x=^1\). So that gives one solution: \(5^{11}=5^1\cdot 5^1\cdot 5^1\cdot 5^1\cdot 5^1\cdot 5^1\cdot 5^1\cdot 5^1\cdot 5^1\cdot 5^1\cdot 5^1\)

OpenStudy (anonymous):

o

OpenStudy (e.mccormick):

\(x=x^1\) missed an x there.

OpenStudy (anonymous):

5^5*5^6

OpenStudy (e.mccormick):

Now, lets take some of those and use the ... yes!

OpenStudy (e.mccormick):

And 6 can be expressed using \((x^{m})^n=x^{m\cdot n}\) as \((5^3)^2\), but I am not sure if they wanted that in this one.

OpenStudy (anonymous):

i dont really think so not yet this is a Quiz it might be on the test

OpenStudy (e.mccormick):

So yah, rules of exponents. If you have items of th same base, like the 5, you can just add up the exponents. \(5^{11}=5^a\cdot 5^b\cdots 5^n\) as long as a to n add up to 11.

OpenStudy (anonymous):

okay

OpenStudy (e.mccormick):

Why \((5^3)^2\) works as a power rule is not too hard if you think about what the outer power means. Multiply everything in the () by itself. So \((5^3)^2=(5^3)(5^3)=5^{3+3}=5^6\)

OpenStudy (e.mccormick):

So there you go, an overview of a ouple exponent rules. Don't have too much fun!

OpenStudy (anonymous):

thank you @e.mccormick

OpenStudy (e.mccormick):

np. Exponents are not too bad, and as you get into them more you will learn how they, roots, and logs are all related.

OpenStudy (anonymous):

\[5^{11}=5^{x+y},where x+y=11\] now you can find the values of x & y satisfying the above eq.